# Improving Text Embeddings with Large Language Models

### 31 Dec 2023 v1

@@ 2 Related Work
Synthetic Data Synthetic data generation is a widely studied topic in information retrieval research, with various methods proposed to enhance retrieval systems with artificially created data. For instance, Doc2query (Nogueira et al., 2019) and Promptagator (Dai et al., 2022) generate synthetic queries for unlabeled documents, which are then leveraged for document expansion or retriever training. GPL (Wang et al., 2022a) employs a cross-encoder to produce pseudo-labels for query-document pairs. Similarly, Query2doc (Wang et al., 2023) generates pseudo-documents for query expansion by few-shot prompting LLMs. Unlike these methods, our approach does not rely on any unlabeled documents or queries and thus can generate more diverse synthetic data.
Another related line of work focuses on knowledge distillation from black-box LLMs by training on synthetic data generated from them. Unnatural Instructions (Honovich et al., 2022) is a synthetic instruction following dataset by prompting existing LLMs. Orca (Mukherjee et al., 2023) and Phi (Gunasekar et al., 2023) propose to train better small language models by using high-quality synthetic data from GPT-3.5 and GPT-4 (OpenAI, 2023).
@@ 5.1 Is Contrastive Pre-training Necessary?
5.2 Personalized Passkey Retrieval
To evaluate the long-context capability of our model, we introduce a novel synthetic task called personalized passkey retrieval, which is illustrated in Figure 4. This task requires encoding the passkey information in a long context into the embeddings. We compare the performance of different variants by changing the sliding window size and the RoPE rotation base (Su et al., 2024) in Figure 5. The results show that the default configuration with k sliding window attains accuracy within k tokens, but the accuracy deteriorates quickly as the context length grows. Naively extending the sliding window size to k results in worse performance. By changing the RoPE rotation base to , the model can achieve over accuracy within k tokens. However, this entails a minor trade-off in performance for shorter contexts. A potential avenue for future research is to efficiently adapt the model to longer contexts through lightweight post-training (Zhu et al., 2023).
@@ Appendix A Implementation Details
Appendix B Instructions for Training and Evaluation

### 19 Jan 2024 v2

@@ 2 Related Work
Synthetic Data Synthetic data generation is a widely studied topic in information retrieval research, with various methods proposed to enhance retrieval systems with artificially created data. For instance, Doc2query (Nogueira et al., 2019), InPars (Bonifacio et al., 2022), and Promptagator (Dai et al., 2022) generate synthetic queries for unlabeled documents, which are then leveraged for document expansion or model training. GPL (Wang et al., 2022a) employs a cross-encoder to produce pseudo-labels for query-document pairs. Similarly, Query2doc (Wang et al., 2023) generates pseudo-documents for query expansion by few-shot prompting LLMs. Unlike these methods, our approach does not rely on any unlabeled documents or queries and thus can generate more diverse synthetic data.
Another related line of work focuses on knowledge distillation from black-box LLMs by training on synthetic data generated from them. DINO (Schick and Schütze, 2021) generates synthetic text pairs for semantic textual similarity. Unnatural Instructions (Honovich et al., 2022) is a synthetic instruction following dataset by prompting existing LLMs. Orca (Mukherjee et al., 2023) and Phi (Gunasekar et al., 2023) propose to train better small language models by using high-quality synthetic data from GPT-3.5/4 (OpenAI, 2023).
@@ 5.1 Is Contrastive Pre-training Necessary?
5.2 Extending to Long Text Embeddings
Existing evaluation datasets for text embedding models are typically short, to evaluate the long-context capability of our model, we introduce a novel synthetic task called personalized passkey retrieval, which is illustrated in Figure 4. This task requires encoding the passkey information in a long context into the embeddings. We compare the performance of different variants by changing the sliding window size and the RoPE rotation base (Su et al., 2024) in Figure 5. The results show that the default configuration with k sliding window attains accuracy within k tokens, but the accuracy deteriorates quickly as the context length grows. Naively extending the sliding window size to k results in worse performance. By changing the RoPE rotation base to , the model can achieve over accuracy within k tokens. However, this entails a minor trade-off in performance for shorter contexts. A potential avenue for future research is to efficiently adapt the model to longer contexts through lightweight post-training (Zhu et al., 2023).
@@ Appendix A Implementation Details
Appendix B Test Set Contamination Analysis
To assess the test set contamination on all the datasets in the MTEB benchmark, we perform a string match based analysis between the test set and our training set, disregarding differences in character case and spacing. We categorize the train-test overlaps into three types:
Low entropy texts. These are texts such as “i need a coffee” and “what does that mean”, which are not considered as contamination because they are common expressions that can occur in various contexts.
Question overlap. We identify test set questions in the DBPedia dataset that also appear in the TriviaQA training set. Given that they constitute a minor portion of the test set, their impact on the overall performance is insignificant.
Retrieval corpus overlap. Several retrieval datasets share the same retrieval corpus. For instance, the DBPedia, NQ, and TriviaQA datasets all use Wikipedia passages, even though their query sets are different. This is a standard evaluation practice in the field of information retrieval, and we do not regard it as contamination.
In summary, we did not detect substantial contamination risks that could alter the main findings of this paper.
Another aspect to consider is the possibility of test set contamination in the training data of Mistral-7B and GPT-4. However, since the training data of these models is not publicly accessible, it is challenging to estimate the degree of such contamination. Given their widespread use in the research community, we believe it is still a valid comparison if other works also employ these models.
@@ Appendix B Test Set Contamination Analysis
Appendix C Prompts for Synthetic Data Generation
For asymmetric tasks, we list the four prompt templates in Table 7, 8, 9, and 10. For symmetric tasks, the prompts templates are available in Table 11 and 12. To generate multilingual data, we sample the value of “{language}” from the language list of XLM-R [Conneau et al., 2020] with higher probability for high-resource languages. When prompting GPT-4/3.5, we set the temperature to and the top- to , which is higher than the default setting to encourage more diversity.
Appendix D Instructions for Training and Evaluation
@@
We manually write instructions for training datasets, as listed in Table 13. For evaluation datasets, the instructions are listed in Table 14.
| Dataset | w/ synthetic only | w/ synthetic + msmarco | w/o synthetic data | full data |
| BIOSSES | 84.2 | 81.0 | 85.4 | 85.5 |
| SICK-R | 78.6 | 78.5 | 81.7 | 82.6 |
| STS12 | 75.8 | 74.7 | 77.9 | 79.7 |
| STS13 | 84.3 | 85.3 | 88.0 | 88.4 |
| STS14 | 80.9 | 81.2 | 83.7 | 84.5 |
| STS15 | 86.2 | 86.8 | 89.5 | 90.4 |
| STS16 | 85.0 | 85.3 | 86.5 | 87.7 |
| STS17 | 87.3 | 87.7 | 91.0 | 91.8 |
| STS22 | 66.0 | 67.1 | 66.2 | 67.0 |
| STSBenchmark | 83.5 | 84.0 | 87.8 | 88.6 |
| SummEval | 31.9 | 32.7 | 31.9 | 31.4 |
| SprintDuplicateQuestions | 93.5 | 95.8 | 96.0 | 95.7 |
| TwitterSemEval2015 | 78.0 | 78.5 | 81.7 | 81.6 |
| TwitterURLCorpus | 86.5 | 86.9 | 87.7 | 87.8 |
| AmazonCounterfactualClass. | 79.6 | 79.9 | 77.2 | 78.7 |
| AmazonPolarityClassification | 95.8 | 95.9 | 93.9 | 95.9 |
| AmazonReviewsClassification | 56.9 | 55.5 | 48.2 | 55.8 |
| Banking77Classification | 86.2 | 87.0 | 88.8 | 88.2 |
| EmotionClassification | 49.2 | 47.6 | 51.0 | 49.8 |
| ImdbClassification | 94.8 | 94.9 | 89.0 | 94.8 |
| MassiveIntentClassification | 79.8 | 79.9 | 79.6 | 80.6 |
| MassiveScenarioClassification | 81.7 | 82.4 | 82.3 | 82.4 |
| MTOPDomainClassification | 95.6 | 95.9 | 95.7 | 96.1 |
| MTOPIntentClassification | 84.9 | 85.9 | 83.4 | 86.1 |
| ToxicConversationsClassification | 70.2 | 70.8 | 70.9 | 69.6 |
| TweetSentimentExtractionClass. | 63.5 | 63.4 | 61.6 | 63.7 |
| AskUbuntuDupQuestions | 64.3 | 65.3 | 67.4 | 67.0 |
| MindSmallReranking | 33.1 | 32.8 | 32.5 | 32.6 |
| SciDocsRR | 86.0 | 86.0 | 85.7 | 86.3 |
| StackOverflowDupQuestions | 52.5 | 53.7 | 55.9 | 54.9 |
| ArxivClusteringP2P | 51.4 | 51.2 | 47.8 | 50.5 |
| ArxivClusteringS2S | 46.5 | 44.9 | 44.6 | 45.5 |
| BiorxivClusteringP2P | 44.5 | 43.3 | 36.9 | 43.5 |
| BiorxivClusteringS2S | 40.9 | 40.1 | 37.0 | 40.2 |
| MedrxivClusteringP2P | 40.5 | 39.9 | 32.6 | 38.2 |
| MedrxivClusteringS2S | 38.0 | 37.9 | 32.8 | 37.5 |
| RedditClustering | 56.3 | 55.9 | 63.1 | 57.7 |
| RedditClusteringP2P | 66.3 | 64.8 | 66.4 | 66.5 |
| StackExchangeClustering | 72.9 | 72.7 | 74.5 | 73.1 |
| StackExchangeClusteringP2P | 46.1 | 45.6 | 34.3 | 45.9 |
| TwentyNewsgroupsClustering | 52.2 | 52.5 | 55.6 | 54.3 |
| ArguAna | 52.2 | 42.7 | 62.5 | 61.9 |
| ClimateFEVER | 21.1 | 28.8 | 25.2 | 38.4 |
| CQADupstackAndroidRetrieval | 40.8 | 36.0 | 44.5 | 43.0 |
| DBPedia | 42.0 | 43.7 | 47.7 | 48.9 |
| FEVER | 72.5 | 83.5 | 73.1 | 87.8 |
| FiQA2018 | 38.1 | 48.4 | 54.5 | 56.6 |
| HotpotQA | 48.1 | 64.0 | 75.6 | 75.7 |
| MSMARCO | 25.7 | 45.0 | 42.9 | 43.1 |
| NFCorpus | 35.5 | 40.0 | 35.3 | 38.6 |
| NQ | 53.3 | 63.5 | 57.3 | 63.5 |
| QuoraRetrieval | 75.0 | 79.5 | 89.5 | 89.6 |
| SCIDOCS | 20.6 | 15.8 | 19.0 | 16.3 |
| SciFact | 71.5 | 71.9 | 74.7 | 76.4 |
| Touche2020 | 25.4 | 32.5 | 19.1 | 26.4 |
| TRECCOVID | 82.3 | 87.3 | 70.8 | 87.2 |
| Average | 63.1 | 64.5 | 64.6 | 66.6 |